# DORSal: Diffusion for Object-centric

### v1

@@ 5.4 Ablations
6 Conclusion
Summary of Limitations.
While DORSal makes significant progress, there are several limitations and open problems worth highlighting, relating to 1) lack of end-to-end training, 2) worse editing performance and consistency for high-resolution training, 3) configuration of the MultiView U-Net architecture for 3D, and 4) non-local editing effects. We discuss these in detail in Appendix B.
Acknowledgments and Disclosure of Funding
Appendix A Broader Impacts
Appendix B Limitations
While DORSal makes significant progress on the challenging problems of novel view synthesis and scene editing, there are several limitations and open problems worth highlighting. As we follow the design of Video Diffusion Models (Ho et al., 2022c) for simplicity, DORSal is not end-to-end trained and is ultimately limited by the quality of the scene representation (Object Slots) provided by the separately trained upstream model (OSRT). End-to-end training comes with additional challenges (e.g. higher memory requirements), but is worth exploring in future work.
As highlighted in our experiments, training at 128x128 resolution with our model design results in decreased editing performance compared to a 64x64 model. We also observed qualitatively worse cross-view consistency in the higher-resolution model. To overcome this limitation, one would likely have to scale the model further in terms of size (at the expense of increased compute and memory requirements) or train a cascade of models to initially predict at 64x64 resolution, followed by one or more conditional upsampling stages, as done in Video Diffusion Models (Ho et al., 2022c).
For scene editing, we find that removing individual object slots can have non-local side effects, e.g. another object or the background changing its appearance, in some cases. Furthermore, edits of individual are typically not perfect, even when trained with a supervised OSRT base model: objects are sometimes only partially removed, or removal of a slot might have no effect at all.
Appendix C Experimental Details
C.1 Evaluation
Novel View Synthesis.
We follow the experimentation protocol outlined in Sajjadi et al. (2022a, c) and evaluate DORSal and baselines using 5 and 3 novel target views for MultiShapeNet and Street View respectively. Similarly, the OSRT base model, is trained with 5 input views on these datasets. To accommodate the U-Net architecture used in DORSal and 3DiM, we crop Street View frames to 128x128 resolution.
Evaluation of Object-level Edits.
We use a median kernel size of 7 for all edit evaluations (incl. the baselines). We evaluate models on the first 1k scenes of the MultiShapeNet dataset. For DORSal, we use an identical initial noise variable (image) for each edit to ensure consistency.
Camera-path Rendering.
For camera-path rendering of many views (beyond what DORSal was trained with) we deploy the sampling procedure outlined in Section 5.3. Camera trajectories follow a circular, center-facing path starting from the first input view. Further, we generate a dataset which has a mix of both nearby views (w.r.t. previously generated views) and uniformly sampled views (at random) from the full view distribution as in MultiShapeNet. Here we train DORSal using 10 input views and 10 target views (down-sampled to 64x64 resolution) to keep a similar amount of diversity when sampling far-away as well as close-by views. 3DiM is trained similarly as for the novel-view synthesis experiments.
C.2 Model Details
C.2.1 DORSal
Conditioning.
We obtain Object Slots from a separately trained OSRT model. In the case of MultiShapeNet, we train OSRT with instance segmentation mask supervision following the approach by Prabhudesai et al. (2023): we take the alpha masks produced by the broadcast decoder to obtain soft segmentation masks, which we match using Hungarian matching with ground-truth instance masks (under an L2 objective) and finally train the model using a cross-entropy loss using the alpha mask logits on the matched target masks. For Street View, we use the default unsupervised OSRT model with a broadcast decoder, as instance masks are not available. All OSRT models use 32 Object Slots.
Network Architecture.
For DORSal we follow the architecture of Ho et al. (2022a), which is a U-Net that has axial attention over time. Differences are that DORSal does not require text conditioning cross attention layers, and it uses slot embeddings augmented with camera poses. The architecture sizes are small compared to Ho et al. (2022a) as can be seen in Table 3. The U-Net on resolutions of 128 $\times$ 128 uses patching to avoid memory expensive feature maps. For the 16 ×$\times 16$ resolution, the ResBlocks use per-view self-attention and between-views cross-attention.
Table 3: DORSal U-Net architecture details.
U-Net 64 192, 384, 576 3 $16\times 16$ No
U-Net 128 256, 512, 1024 3 $16\times 16$ $2\times 2$
Training.
We adopt a similar training set-up to Ho et al. (2022c), using a cosine-shaped noise schedule, a learning rate with a peak value of 0.00003 using linear warm-up for 5000 steps, optimization using Adam with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and EMA decay for the model parameters. We train with a global batch size of 8 and classifier-free guidance with a conditioning dropout probability of $0.1$ (and an inference guidance weight of $2$). We report results after training for 1 000 000 steps. For MultiShapeNet, we use Object Slots from Sup-OSRT (i.e. supervised) and do not use slot dropout during training. For Street View, we use Object Slots from OSRT (i.e. unsupervised) and use a slot dropout probability of 0.2, which we found to improve editing quality on this dataset (compared to no slot dropout).
Camera-Path Sampling.
We leverage the iterative nature of the generative denoising process to create smooth transitions as well as global consistency between frames. Our technique (Figure 9) is inspired by Hoogeboom et al. (2023), where high resolution images are generated with overlapping patches by dividing the typical denoising process with $T$ steps is divided into multiple stages. Instead of generating a patch entirely for all denoising timesteps from one to zero, first a noisy version of each patch is generated by generating from one to $3/4$, in the case of 4 stages. Only after all noisy patches at time $3/4$ are generated, do they proceed to the next stage from $3/4$ to $2/4$.
To apply this technique for 3D camera-path rendering of up to 190 views in Section 5.3, we propose to interleave 3 types of frame shuffling for subsequent stages: 1) no shuffle (identity), to allow the model to make blocks of the context length consistent; 2) shift the frames in time by about half of the context length, which puts frames together with new neighbours in their context, allowing the model to create smooth transitions; 3) shuffle all frames with a random permutation, to allow the model to resolve inconsistencies globally.
C.2.2 3D Diffusion Model (3DiM)
We compare to 3DiM, which is a pose-conditional image-to-image diffusion model for generating novel views of the same scene (Watson et al., 2023). During training, 3DiM takes as input a pair of views of a static scene (including their poses), where one of the views (designated as the “target view”) is corrupted with noise. The training objective is to predict the Gaussian noise that was used to corrupt the target view. During inference, 3DiM makes use of stochastic conditioning to generate 3D-consistent views of a scene. In particular, given a small set of $k$ conditioning views and their camera poses (typically $k=1$), a new view for a given target camera pose is generated by conditioning on a randomly selected view from the conditioning set at each denoising step. Each time a new view is generated, it is added to the conditioning set.
Network Architecture.
[…]
Training.
We adopt the same training set-up as in the 3DiM paper, which consist of a cosine-shaped noise schedule, a learning rate with peak value of 0.0001 using linear warm-up for 10M samples, optimization using Adam with $\beta_{1}=0.9$ and $\beta_{2}=0.99$, and EMA decay for the model parameters. We train with a global batch size of 128 and classifier-free guidance $10\%$ with a weight of $3$, as was done for the experiment on SRN cars in their paper. We report results after training for 320 000 steps.
Sampling.
We generate samples in the same way as in the 3DiM paper, using 256 DDPM denoising steps and clip to [-1, 1] after each step.
For additional details, including code, we refer to Sections 6 & 7 in Watson et al. (2023).
C.2.3 SRT & OSRT
SRT was originally proposed by Sajjadi et al. (2022c) with Set-Latent Scene Representations (SLSR) and subsequently adapted to Object Slots for OSRT Sajjadi et al. (2022a). At the same time, a few tweaks were made to the model, e.g. by using a smaller patch size and a larger render MLP Sajjadi et al. (2022a). For all our experiments (SRT and OSRT), we use the improved architecture from the OSRT paper, see Appendix A.4 in Sajjadi et al. (2022a) for the model and training details. Following Sajjadi et al. (2022c), we train all models for $\sim$4M steps on both datasets.
C.3 Compute and Data Licenses
Appendix D Additional Results
D.1 Qualitative Results
@@
We provide additional qualitative results for novel view synthesis in Figure 10 (Street View) and in Figure 11 (MultiShapeNet). For Street View, it is evident that even when modifying 3DiM to use 5 ground-truth input views during inference, it is unable to synthesize accurate views from novel directions, while DORSal renders realistic views that adhere to the content of the scene.
@@
In Figure 12 we provide exhaustive scene editing results for several Street View scenes: each image shows one generation of DORSal with exactly one slot removed. These results further highlight that several meaningful edits can be made per scene. Typical failure modes can also be observed: 1) some objects are unaffected by slot removal, 2) some edits have side effects (e.g. another object disappearing or changing its appearance), and 3) multiple different edits have the same (or a very similar) effect. These failure modes likely originate in part from the unsupervised nature of the OSRT base model, which sometimes assigns multiple slots to a single object, or does not decompose the scene well. Fully “imagined” objects (i.e. objects which are not visible in the input views and therefore not encoded in the Object Slots) further generally cannot be directly edited in this way. Some of these issues can likely be overcome in future work by incorporating object supervision (as done for MultiShapeNet), and by devising a mechanism by which “imagined” objects not visible in input views are similarly encoded in Object Slots.
Figure 12: Scene Editing (Street View). Exhaustive DORSal scene editing results for three Street View scenes, with one Object Slot removed at a time. Several examples where scene content differs are highlighted.
Table 4: Novel-View Synthesis. Evaluation on 1 k test scenes from MultiShapeNet and Street View.
@@
Model PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$ FIDnormal-$\downarrow$ PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$ FIDnormal-$\downarrow$
3DiM 18.20 0.559 0.287 10.94 12.68 0.283 0.477 15.58
3DiM (5 input) 21.46 0.707 0.182 8.20 12.25 0.271 0.557 34.47
DORSal 18.76 0.557 0.266 11.01 16.05 0.421 0.361 16.24
D.2 Sampling from DORSal using DDIM vs. DDPM
In Table 5 we additionally report results when sampling from DORSal using 256 steps of DDPM. The latter lends itself better when using longer sampling chains as is reflected in an improvement in FID and PSNR (Song et al., 2021). Notice how in this case, when using the same procedure for sampling, DORSal consistently outperforms 3DiM.
Table 5: Novel-View Synthesis (DDIM vs. DDPM). Evaluation on 1 k test scenes from MultiShapeNet and Street View.
MultiShapeNet Street View
Model PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$ FIDnormal-$\downarrow$ PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$ FIDnormal-$\downarrow$
3DiM 18.20 0.559 0.287 10.94 12.68 0.283 0.477 15.58
DORSal 18.76 0.557 0.266 11.01 16.05 0.421 0.361 16.24
DORSal (DDPM) 18.99 0.557 0.265 9.00 16.36 0.425 0.356 14.62
D.3 Comparison to 3DiM using Additional Input Views
@@
D.4 Full Edit Metrics (incl. Background Slots)
Table 6 presents quantitative results for scene editing when additionally including the background slots during evaluation. For Edit-mIoU and Edit-ARI effectively the same conclusions hold true: DORSal performs well compared to OSRT and attains performance close to that of Sup-OSRT. Further, it can be observed how the smaller DORSal model is more capable of scene editing compared to the large model trained on higher resolution data. A noticeable difference to the results reported in Table 2 can be observed for OSRT, where we see a large drop in performance when reporting Edit-ARI (compared to Edit-FG-ARI). This difference suggests that the object segments OSRT learns only loosely correspond to the precise object shape, but rather spans a part of the background as well. As a consequence, modifying individual slots will also significantly alter the background pixels in an undesirable way.
Table 6: Scene Editing. Evaluation on 1k test scenes of MultiShapeNet. All metrics reported in %.
Model Edit-mIoU Edit-FG-mIoU Edit-ARI Edit-FG-ARI
OSRT 43.4 43.1 38.2 79.6
Sup-OSRT 52.0 50.0 79.4 75.5
DORSal (64x64) 47.4 45.8 67.9 70.0
DORSal (128x128) 36.6 35.0 53.5 64.6

### v3

@@ C.1 Ablations
C.2 Variance
While we report results for a single representative model run in the main paper, we found that variance in quantitative metrics across different runs is fairly moderate. Specifically, we measure a standard error of approx. 0.004 for LPIPS and approx. 0.2 for PSNR (for 3 model training re-runs with different initialization seeds), which does not affect the interpretation of our reported results.
C.3 3D Consistency
To provide some further insight into the 3D consistency of DORSal, we re-computed the edit-segmentation scores on a subset of the samples where we measure both Edit FG-ARI across all the views simultaneously and a “2D Edit FG-ARI” where we compute Edit FG-ARI for each view individually and then average the results. Note that the latter does not penalize inconsistencies between the views, hence the gap between the two scores is indicative of any inconsistencies taking place. A similar approach to evaluating 3D consistency was carried out in Sajjadi et al. (2022a). We obtain 0.702 Edit FG-ARI and 0.721 2D Edit FG-ARI. The small gap between these scores indicates that the segmentation obtained via this procedure is highly consistent across views.
C.4 Reproducing Input Views
Table 5: Reproducing input views. Eval on 16 MultiShapeNet scenes.
Model PSNR LPIPS
Sup-OSRT 23.72 0.340
DORSal 19.55 0.252
To give an indication of whether the slots contain sufficient information about the scene, we ran DORSal and sup-OSRT on 16 scenes of MultiShapeNet with target views matching input views (Table 5). For both models, we observe how predicting input views improves PSNR and LPIPS (compared to their respective novel-view synthesis scores in Table 1), indicating that the models are better able to leverage the information contained in object slots in this setting. In particular, the significantly lower LPIPS score for DORSal indicates faithful reconstruction of the inputs. This is a positive result for DORSal in particular, suggesting that it hallucinates less in these instances.
C.5 Qualitative Results
@@
We provide additional qualitative novel view synthesis results in Figures 10 (Street View) and 11 (MultiShapeNet). For Street View, it is evident that even when modifying 3DiM to use 5 ground-truth input views during inference, it is unable to synthesize accurate views from novel directions, while DORSal renders realistic views that adhere to the content of the scene.
@@
We show qualitative results for scene editing in MultiShapeNet in Figure 12 and for Street View in Figure 13. In Figure 14 we further provide exhaustive scene editing results for several Street View scenes: each image shows one generation of DORSal with exactly one slot removed. These results further highlight that several meaningful edits can be made per scene. Typical failure modes can also be observed: 1) some objects are unaffected by slot removal, 2) some edits have side effects (e.g. another object disappearing or changing its appearance), and 3) multiple different edits have the same (or a very similar) effect. These failure modes likely originate in part from the unsupervised nature of the OSRT base model, which sometimes assigns multiple slots to a single object, or does not decompose the scene well. Fully “imagined” objects (i.e. objects which are not visible in the input views and therefore not encoded in the Object Slots) further generally cannot be directly edited in this way. Some of these issues can likely be overcome in future work by incorporating object supervision (as done for MultiShapeNet), and by devising a mechanism by which “imagined” objects not visible in input views are similarly encoded in Object Slots.
Figure 12: Scene Editing (MultiShapeNet). We remove one slot at a time in the conditioning of DORSal and render the resulting scene while keeping the initial image noise fixed. In the leftmost panel, the slot corresponding to the background is removed while all objects are present. The other panels show deleted objects (highlighted in red circles) when their corresponding slot is removed.
Figure 13: Scene Editing (Stree View). Removing one slot at a time, we here show further examples on the Street View dataset where objects are erased from the scene.
Figure 14: Scene Editing (Street View; Exhaustive). Exhaustive DORSal scene editing results for three Street View scenes, with one Object Slot removed at a time. Several examples where scene content differs are highlighted.
Table 6: Novel-view synthesis. Using additional ground-truth input views for 3DiM.
@@
[…]
3DiM 18.20 0.287 10.94 12.68 0.477 15.58
3DiM (5 input) 21.46 0.182 8.20 12.25 0.557 34.47
DORSal (DDPM) 18.99 0.265 9.00 16.36 0.356 14.62
C.6 Comparison to 3DiM using Additional Input Views